OpenStudy (anonymous):

A rectangle has a length 2 ft less than three times its width. What is the area of the rectangle when the width is 2 ft?

OpenStudy (nincompoop):

read the sentence that has a question mark at the end

OpenStudy (anonymous):

@nincompoop can you help me?

OpenStudy (nincompoop):

exactly what I am doing now read the the sentence that has a question mark at the end

OpenStudy (anonymous):

omg its 2ft isn't it

OpenStudy (nincompoop):

?

OpenStudy (nincompoop):

sentence, meg

OpenStudy (anonymous):

what about it?

OpenStudy (nincompoop):

what was the question it was asking?

OpenStudy (anonymous):

"what is the area of the rectangle when the width is 2ft?"

OpenStudy (nincompoop):

What is the area of the rectangle when the width is 2 ft? AREA OF RECTANGLE

OpenStudy (nincompoop):

do you know how to solve for an area of a rectangle?

OpenStudy (anonymous):

yes its A=L x W correct?

OpenStudy (nincompoop):

good

OpenStudy (nincompoop):

what is the value of the width?

OpenStudy (anonymous):

4ft?

OpenStudy (nincompoop):

where do you see 4? What is the area of the rectangle when the width is 2 ft? What is the area of the rectangle when the width is 2 ft? What is the area of the rectangle when the width is 2 ft?

OpenStudy (nincompoop):

so what is the value of the width?

OpenStudy (anonymous):

8ft?

OpenStudy (nincompoop):

OMG the value of the width is in the question What is the area of the rectangle when the \(\huge width~is 2~ft?\)

OpenStudy (nincompoop):

so what is the value of the width?

OpenStudy (anonymous):

2

OpenStudy (nincompoop):

haha good!

OpenStudy (anonymous):

but 2 isn't one of the answers

OpenStudy (nincompoop):

now rewrite the area formula like you did before but now replace W with 2

OpenStudy (anonymous):

OH

OpenStudy (anonymous):

okay A=L x 2

OpenStudy (nincompoop):

good!

OpenStudy (nincompoop):

what do we have to figure out next?

OpenStudy (anonymous):

the length

OpenStudy (shadowlegendx):

Slow and steady

OpenStudy (nincompoop):

good, and what is our clue for its value?

OpenStudy (shadowlegendx):

;)

OpenStudy (nincompoop):

A rectangle has a length 2 ft less than three times its width.

OpenStudy (anonymous):

sorry its 4 am here my brain isn't quite awake and our clue is 2ft because it it says

OpenStudy (nincompoop):

2 ft less than 3 times its width JUST BEAR WITH ME :) we will be done soon and you would be glad you stayed the course

OpenStudy (anonymous):

2ft-3ft?

OpenStudy (nincompoop):

it says 2 less than 3 times its width we know that width is 2 3 times its with is written as \(\color{red}{3 \times 2} \) am I correct or am I correct?

OpenStudy (anonymous):

correct

OpenStudy (nincompoop):

and it says 2 less than (2 x 3)

OpenStudy (nincompoop):

so it means it is 2 units shorter than the 6

OpenStudy (nincompoop):

what is 2 subtracted out of 6?

OpenStudy (anonymous):

where did you get 6 from?

OpenStudy (anonymous):

4ft

OpenStudy (nincompoop):

3 times its width <-----------

OpenStudy (nincompoop):

what is the value of the width again?

OpenStudy (anonymous):

6ft

OpenStudy (nincompoop):

no! the value of width is 2

OpenStudy (nincompoop):

remember?

OpenStudy (nincompoop):

What is the area of the rectangle when the width is 2 ft? remember this part?

OpenStudy (anonymous):

oh yes!!! good god I'm not a smart person

OpenStudy (anonymous):

yes yes

OpenStudy (nincompoop):

you are not smart, but you can start with paying attention to the details you are given and work with those

OpenStudy (nincompoop):

so 3 times 2 is?

OpenStudy (anonymous):

6ft

OpenStudy (nincompoop):

great!

OpenStudy (nincompoop):

but the problem says A rectangle has a length 2 ft less than three times its width so what is 2 ft less than 6 ft?

OpenStudy (anonymous):

4ft

OpenStudy (nincompoop):

you're basically subtracting 2ft out of 6 ft

OpenStudy (nincompoop):

GREAT! we are making progress

OpenStudy (anonymous):

good god I'm sorry for wasting your time :((

OpenStudy (nincompoop):

rewrite the formula of rectangle area again but use the numerical value and their units

OpenStudy (nincompoop):

are you doing it?

OpenStudy (anonymous):

so A= L x W 6= 2-3?

OpenStudy (anonymous):

no thats not right,

OpenStudy (nincompoop):

yes, that is not right

OpenStudy (nincompoop):

we are looking for A we only found the values of L and W so far

OpenStudy (anonymous):

A=L x W A= 2ft x 3ft

OpenStudy (nincompoop):

GREAT! AWESOME!

OpenStudy (anonymous):

YES!

OpenStudy (nincompoop):

pay close attention okay? let us replace ft into a single letter, say the letter \(\huge a \) deal?

OpenStudy (anonymous):

ok

OpenStudy (nincompoop):

by doing that, we have \(\sf A = 4a \times 2a \)

OpenStudy (nincompoop):

can you try to multiply those for me?

OpenStudy (anonymous):

alright.. is 8a

OpenStudy (nincompoop):

you got the coefficient right, but not the variable \(\large a \)

OpenStudy (nincompoop):

so it is clearer that can be written also as \(\color{red}{4\times 2 \times a \times a} \)

OpenStudy (nincompoop):

what is a x a ?

OpenStudy (anonymous):

OH OF COURSE so 8^2?

OpenStudy (nincompoop):

do not forget the a

OpenStudy (anonymous):

8^2a?

OpenStudy (nincompoop):

laughing hysterically inside my head the squared is on a not on 8

OpenStudy (nincompoop):

try again

OpenStudy (nincompoop):

\(\color{red}{4\times 2 \times a \times a}\)

OpenStudy (anonymous):

okay.. 8a^2?

OpenStudy (nincompoop):

YESSSSSSS! omg you did it!

OpenStudy (nincompoop):

remember that we replaced ft into a?

OpenStudy (anonymous):

yes

OpenStudy (nincompoop):

here's the reason why I asked you to do that I want to make sure that you follow the same rules for units (ft, meters, ounce, etc) as when you're multiplying variables

OpenStudy (anonymous):

makes sense :)

OpenStudy (nincompoop):

now rewrite your \(8a^2 \) into proper unit

OpenStudy (anonymous):

so the answer is 8ft^2?

OpenStudy (nincompoop):

YES!

OpenStudy (anonymous):

YAAAAS

OpenStudy (nincompoop):

now, tell me if you regret staying or you're glad you stayed the course?

OpenStudy (anonymous):

i regret absolutely nothing I'm glad i stayed

OpenStudy (nincompoop):

thank you for being awesome and have a nice rest

OpenStudy (anonymous):

thanks for spending time on a person like me nincompoop :)

HanAkoSolo (jamierox4ev3r):

@nincompoop I applaud you tremendously. Outstanding work